OpenStudy (anonymous):

Jackson goes to the gym 0, 2, or 3 days per week, depending on work demands. The expected value of the number of days per week that Jackson goes to the gym is 2.05. The probability that he goes 0 days is 0.1, the probability that he goes 2 days is_____ , and the probability that he goes 3 days is____ .

OpenStudy (anonymous):

@iGreen @Blonde_Gangsta

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

I am not current 7 days in one week 0*0.01 + 2x + 3 y = 2.05 x + y = 7 consider how many outcome for 2 days out of 7? how many ways 3 days out of 7 let me try and solicit some help

OpenStudy (triciaal):

@Michele_Laino will you please help

OpenStudy (anonymous):

ok thanks

OpenStudy (michele_laino):

I think that the requested probability can be found, solving the subsequent system: \[\Large \left\{ \begin{gathered} 0 \times 0.1 + x \times 2 + y \times 3 = 2.05 \hfill \\ x + y + 0 = 1 \hfill \\ \end{gathered} \right.\] where x, and y represent the requested probability. In particular, the second equation means that the sum of all probabilities, has to be equal to 1 Whereas the first equation, comes from the definition of the mean value

OpenStudy (anonymous):

ok...@Michele_Laino

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

what values for x, and for y, do you get?

OpenStudy (anonymous):

@Michele_Laino the \(0\) in the second equation should be \(0.1\).

OpenStudy (anonymous):

@SithsAndGiggles can you tell me how to do this?

OpenStudy (anonymous):

@MrNood can you tell me how to do it?

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@amistre64 @jdoe0001

OpenStudy (amistre64):

\[E(X)=\sum x_iP(x_i)\]

OpenStudy (amistre64):

\[\sum P(x_i)=1\]

OpenStudy (amistre64):

x = 0 2 3 P(x) = 0.1 a b 0(.1) + 2a + 3b = 2.05 0.1 + a + b = 1

OpenStudy (anonymous):

how would you do the equation? @amistre64

OpenStudy (amistre64):

you have 2 equations in 2 unknowns, a very basic algebra setup. Algebra is usually a class taken before stats.